# DialSim: A Real-Time Simulator for Evaluating Long-Term Multi-Party Dialogue Understanding of Conversation Systems

### v5

@@ Appendix I Experimental Results in an Unlimited Time Setting
Table 14: The performances of the agents on Friends dialogue in DialSim (time limit = 6 seconds, with shuffled names). We conducted experiments three times and reported the accuracies and the standard deviations.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 32.18 (1.74) 30.27 (0.73) 38.40 (0.30) 34.70 (0.22) 32.90 (0.57) 36.10 (0.36) 36.40 (0.73) Open Tülu 2-70B 2.54 (0.21) 26.47 (1.91) 31.75 (1.71) 30.94 (2.41) 31.90 (1.25) 29.83 (1.03) 31.86 (2.13) Tülu 2-7B 1.15 (0.06) 28.20 (1.63) 27.64 (2.37) 27.78 (1.32) 28.98 (0.96) 25.03 (0.94) 29.08 (2.47) Llama3.1-70B 31.38 (1.01) 29.08 (1.57) 36.48 (2.51) 36.91 (0.36) 35.89 (0.65) 39.80 (1.42) 39.00 (0.87) Llama3.1-8B 27.16 (1.62) 25.76 (1.42) 30.61 (1.25) 29.59 (1.25) 30.91 (0.99) 29.76 (1.26) 31.59 (0.69) Mixtral-8x7B 34.19 (0.68) 25.23 (1.19) 37.72 (0.96) 29.48 (0.87) 29.09 (1.46) 31.78 (1.71) 29.45 (0.04) Mistral-7B 27.78 (1.62) 25.02 (1.26) 30.65 (1.39) 24.99 (1.51) 27.34 (0.49) 27.97 (1.31) 26.97 (1.45) Gemma-7B 17.98 (2.15) 21.64 (0.39) 22.31 (2.15) 18.66 (1.55) 25.97 (1.92) 21.79 (0.40) 21.22 (0.59) Gemma-2B 1.04 (0.19) 24.19 (0.82) 25.25 (1.02) 24.32 (1.55) 25.03 (0.66) 25.44 (1.96) 23.62 (0.36)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
@@
Table 15: The performances of the agents on Friends dialogue in DialSim (without a time limit and with shuffled names). We conducted experiments three times and reported the accuracies and the standard deviations.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 31.72 (1.69) 29.29 (0.74) 42.95 (2.04) 31.12 (0.94) 35.08 (1.02) 38.14 (3.63) 36.91 (2.00) Open Tülu 2-70B 0.21 (0.07) 18.24 (0.84) 20.60 (1.00) 18.64 (1.81) 31.71 (2.22) 7.82 (1.57) 17.31 (0.61) Tülu 2-7B 0.74 (0.15) 13.19 (0.35) 19.54 (1.29) 26.07 (2.00) 13.87 (0.71) 18.35 (1.21) 27.48 (2.04) Llama3.1-70B 0.64 (0.10) 29.29 (0.59) 0.60 (0.12) 15.07 (5.12) 39.08 (0.99) 2.43 (0.10) 18.18 (0.16) Llama3.1-8B 26.61 (1.24) 26.86 (0.78) 31.20 (1.87) 27.08 (0.63) 24.82 (1.07) 31.72 (1.66) 22.69 (0.63) Mixtral-8x7B 2.41 (0.40) 14.90 (0.82) 23.55 (0.40) 15.64 (0.47) 16.43 (1.68) 22.95 (0.68) 13.22 (1.61) Mistral-7B 3.35 (0.58) 24.44 (1.13) 31.39 (0.70) 24.26 (1.60) 29.82 (0.95) 30.21 (0.90) 23.90 (0.51) Gemma-7B 18.05 (0.97) 22.52 (0.81) 20.64 (0.26) 16.63 (1.59) 23.41 (1.26) 18.34 (0.82) 19.48 (2.45) Gemma-2B 0.47 (0.13) 24.31 (0.96) 24.77 (0.74) 25.74 (1.46) 28.41 (1.20) 24.68 (1.45) 24.75 (1.50)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
Table 16: The performances of the agents on Friends dialogue in DialSim (time limit = 6 seconds, with new names replaced). We conducted experiments three times and reported the accuracies and the standard deviations.
Type Model Base LLM RAG-based BM25 OpenAI Embedding Utterance Session Entire Session Sum. Utterance Session Entire Session Sum. API GPT-4o-mini 32.06 (0.52) 30.01 (1.09) 46.57 (1.56) 34.44 (1.53) 34.23 (0.99) 38.78 (0.24) 37.97 (1.24) Open Tülu 2-70B 2.17 (0.46) 27.24 (1.17) 33.34 (1.17) 32.85 (1.85) 34.95 (0.47) 29.41 (1.22) 33.55 (2.79) Tülu 2-7B 0.63 (0.26) 30.26 (1.03) 27.68 (1.24) 30.98 (1.08) 30.99 (0.22) 27.93 (1.97) 31.80 (2.05) Llama3.1-70B 31.03 (1.91) 28.91 (2.33) 38.44 (5.98) 41.68 (3.68) 38.40 (1.10) 40.83 (1.07) 44.27 (0.57) Llama3.1-8B 26.65 (1.19) 25.80 (0.18) 32.01 (1.10) 30.48 (1.30) 29.50 (1.10) 32.82 (0.63) 32.23 (2.32) Mixtral-8x7B 38.92 (1.61) 26.91 (1.46) 39.98 (2.98) 27.02 (0.42) 30.27 (1.37) 35.89 (0.37) 29.52 (1.28) Mistral-7B 29.10 (1.34) 23.33 (0.83) 34.59 (0.80) 27.87 (2.66) 30.59 (2.09) 30.45 (0.89) 27.93 (0.99) Gemma-7B 17.37 (0.77) 22.58 (1.62) 21.41 (1.53) 21.61 (1.53) 23.90 (0.90) 21.61 (1.09) 20.88 (0.91) Gemma-2B 0.37 (0.07) 23.48 (1.85) 23.13 (1.14) 25.72 (2.67) 29.12 (1.90) 24.88 (1.40) 24.81 (1.31)
$\dagger$: Both GPT-4o-mini and Llama3.1 support up to 128k tokens, but we limited them to 8k tokens due to high costs and GPU VRAM limits, respectively.
Table 17: The performances of the agents on Friends dialogue in DialSim (without a time limit and with new names replaced). We conducted experiments three times and reported the accuracies and the standard deviations.

### v6

@@ Appendix I Experimental Results in the Oracle Setting
Table 13: The performance of the agents on original Friends dialogue in DialSim. We conducted experiments three times and reported the accuracy and standard deviations. Bold indicates the best performance per retrieval method.
Type Model Base LLM RAG-based
BM25 OpenAI Embedding
Utterance Session Entire Session Sum. Utterance Session Entire Session Sum.
API GPT-4o-mini 50.28 (0.51) 44.00 (2.62) 55.62 (0.45) 48.66 (3.58) 47.45 (0.45) 52.81 (1.21) 50.70 (0.64)
GPT-4.1-nano 22.61 (1.31) 24.22 (1.52) 29.63 (1.40) 26.18 (0.89) 30.14 (0.26) 26.52 (1.62) 27.03 (1.05)
Gemini 2.5 Flash 60.34 (6.32) 33.08 (0.77) 56.26 (1.47) 25.71 (2.21) 40.10 (0.13) 50.19 (0.38) 26.63 (0.06)
Gemini 2.0 Flash 44.91 (1.81) 32.44 (0.68) 54.34 (2.11) 25.42 (0.51) 36.97 (0.45) 45.72 (1.40) 26.12 (1.72)
Open Llama 3.3-70B 36.06 (2.03) 35.89 (3.18) 48.23 (1.04) 45.04 (2.61) 41.85 (2.72) 43.51 (1.51) 49.00 (2.47)
Llama 3.1-8B 26.91 (1.12) 29.89 (1.56) 34.70 (1.75) 33.93 (1.76) 31.63 (2.17) 32.91 (0.51) 35.59 (1.09)
Mixtral-8x7B 42.19 (1.76) 31.84 (0.78) 46.47 (1.75) 32.31 (1.09) 35.51 (0.19) 41.24 (2.90) 34.18 (0.96)
Mistral-7B 32.93 (0.59) 28.20 (1.17) 35.09 (1.76) 30.16 (1.82) 30.12 (1.45) 31.00 (1.93) 30.80 (1.75)
Qwen 3-32B 36.10 (1.16) 33.59 (1.16) 51.72 (2.68) 44.13 (0.70) 38.57 (0.21) 45.72 (0.91) 45.00 (0.30)
Qwen 3-8B 28.86 (1.64) 29.25 (1.28) 39.72 (1.13) 34.61 (1.91) 33.63 (1.04) 37.59 (0.95) 35.25 (1.61)
Phi 4-14B 24.18 (0.84) 27.12 (1.84) 37.93 (2.63) 29.93 (1.24) 30.61 (1.68) 31.46 (1.26) 33.63 (1.48)
Phi 4 mini-3.8B 10.34 (0.81) 25.71 (1.22) 30.18 (1.24) 20.99 (0.57) 30.99 (2.83) 24.69 (0.87) 21.16 (0.06)
Random Guessing 20.00 20.00 20.00 20.00 20.00 20.00 20.00
@@
Table 14: The performance of the agents on the adversarial version of Friends dialogue in DialSim. We conducted experiments three times and reported the accuracy and standard deviations. Bold indicates the best performance per retrieval method.
Type Model Base LLM RAG-based
BM25 OpenAI Embedding
Utterance Session Entire Session Sum. Utterance Session Entire Session Sum.
API GPT-4o-mini 39.72 (0.75) 33.84 (0.77) 45.08 (1.92) 37.36 (1.09) 36.61 (2.14) 41.38 (0.54) 42.10 (1.04)
GPT-4.1-nano 23.56 (1.34) 25.22 (0.70) 27.20 (0.51) 24.07 (0.70) 24.78 (1.18) 27.54 (1.99) 24.90 (0.68)
Gemini 2.5 Flash 58.81 (3.90) 28.57 (0.42) 51.40 (0.57) 16.77 (9.72) 34.99 (0.77) 43.23 (0.19) 24.78 (0.64)
Gemini 2.0 Flash 36.74 (3.66) 28.14 (0.97) 47.81 (0.39) 23.54 (1.10) 33.80 (0.91) 38.53 (0.81) 22.09 (0.75)
Open Llama 3.3-70B 31.38 (1.01) 29.50 (1.13) 42.87 (1.12) 37.85 (1.22) 35.67 (0.57) 38.61 (2.28) 39.85 (0.63)
Llama 3.1-8B 26.99 (2.56) 27.25 (1.87) 38.48 (2.41) 30.65 (1.01) 29.03 (1.34) 38.57 (0.81) 32.99 (1.42)
Mixtral-8x7B 34.19 (0.68) 25.23 (1.19) 37.72 (0.96) 29.48 (0.87) 24.27 (1.40) 31.80 (0.38) 25.61 (1.21)
Mistral-7B 26.44 (1.68) 25.24 (1.55) 33.33 (1.26) 26.31 (0.63) 27.25 (0.59) 29.76 (1.31) 28.95 (1.51)
Qwen 3-32B 32.31 (0.73) 30.86 (1.42) 46.32 (2.40) 36.31 (1.39) 33.89 (0.47) 39.04 (0.43) 37.42 (0.38)
Qwen 3-8B 27.08 (0.65) 27.08 (1.15) 38.95 (0.48) 30.95 (1.72) 32.78 (1.68) 34.18 (1.56) 34.14 (2.68)
Phi 4-14B 19.41 (1.01) 26.35 (0.47) 33.04 (1.20) 26.18 (1.18) 27.29 (0.59) 29.63 (0.99) 31.08 (2.21)
Phi 4 mini-3.8B 8.98 (0.42) 26.01 (0.43) 30.82 (0.34) 21.75 (0.16) 31.59 (0.69) 26.14 (2.07) 23.33 (0.69)
Random Guessing 20.00 20.00 20.00 20.00 20.00 20.00 20.00